OpenStudy (anonymous):

A rope, under a tension of 272 N and fixed at both ends, oscillates in a second-harmonic standing wave pattern. The displacement of the rope is given by y = 0.40sin(πx/2.0) sin(13πt), where x = 0 at one end of the rope, x and y are in meters, and t is in seconds. The length of the rope is A.1 m B. 2 m C. 4 m D. 8 m E. 16 m The mass of the rope is in the previous question is A. 1.61 kg B. 1.92 kg C. 2.33 kg D. 5.14 kg E. 8.97 kg

OpenStudy (anonymous):

i get 4m for the length of the rope PI/2=(2*PI)/landa =4m but thats wrong!

OpenStudy (anonymous):

oh just noticed its a standing wave...

OpenStudy (anonymous):

does that mean i multiply k by 2?

OpenStudy (anonymous):

so then landa equals 2 2(PI/2)=(2*PI)/landa PI*landa=2*PI landa=2

OpenStudy (anonymous):

lost for the next bit

OpenStudy (anonymous):

ok i just realised i have 2 practice papers and one says 4m and the other says 2m. now i am really lost lol

OpenStudy (anonymous):

if its second harmonic L=landa=(2*landa)/(landa) so it still is landa so i was right the first time...its 4m. so one past paper is right. am i right?

OpenStudy (vincent-lyon.fr):

Length is 4 metres and mass is 1.61 kg.

OpenStudy (anonymous):

ah k so i just realised that \[\omega=\sqrt{k/m}\] but i don't understand the next bit \[m=T(k/\omega)^{2}L\] what is the L? that was in my lecture notes

OpenStudy (anonymous):

it is supposed to be 4

OpenStudy (anonymous):

can you show working and explanation of formulas used?

OpenStudy (anonymous):

please

OpenStudy (aravindg):

hmmm...

OpenStudy (anonymous):

hmmm indeed

OpenStudy (anonymous):

its a conspiracy

OpenStudy (vincent-lyon.fr):

If you have second-harmonic, amplitude must be zero at x=0, x=L/2 and x=L. Given sin(πx/2.0), this implies L=4.0m so \(\lambda\)=4.0m and \(\omega\)=13\(\pi\) Hence c=\(\lambda \nu=\Large\frac{\lambda\omega}{2\pi}\)=26 m/s For a string: c=\(\Large \sqrt \frac T\mu\) so \(\mu=\frac {T}{c^2}\)=0.40 kg/m M = Lµ = 1.6 kg